# A Novel Framework: From Trace Representation to Abnormal Trace Classification in Microservice-based systems

### v1

@@ III-C2 Meta-training
The standard MAML outer loop update involves a computationally intensive process of computing a gradient through a gradient, which requests an additional backward pass through the function $f$ to compute Hessian-vector products [finn2017model]. We use a first-order approximation [finn2017model] to simplify the update by such second-order derivatives.
@@ IV-A1 Dataset
We use two open-source datasets to establish our own dataset: DeepTraLog [DeepTraLog] and Nezha [Nezha]. (1) DeepTraLog was collected from TrainTicket, a medium-scale, widely-used benchmark MSS for train ticket booking, comprising 45 services. TrainTicket contains various real-industrial faults in different branches. DeepTraLog was created by Zhang et al. [zhang2022deeptralog], who executed such fault branches of TrainTicket and generated abnormal traces of 14 fault categories. These fault categories pertain to related to asynchronous interaction, multi-instance, configuration, and monolithic dimensions. (2) Nezha was created by Yu et al. [yu2023nezha]. It includes data from two MSS: TrainTicket, and a small-scale e-commerce benchmark MSS OnlineBoutique composed of 12 services. This dataset was constructed by injecting faults into the pods of TrainTicket and OnlineBoutique and collocating abnormal traces. Within each system, faults injected into each pod belong to four categories: CPU-intention, CPU-consumption, service exception, and message return. Each fault case within a pod here is considered as a unique fault category.
@@ IV-A2 Training and evaluation
E1 and E3 are within system experiments while E2 and E4 are cross-system experiments. Considering prior MAML studies [ye2021train, finn2017model] and our fault dataset, we investigate 5-way 5-shot and 10-shot setups across experiments E1-E4. In these experiments, each meta-training and meta-testing task is a unique abnormal trace classification task, where each task involves classifying abnormal traces from a distinct set of 5-way fault categories. In each experiment, we train and evaluate our TE-MAML using the randomly established meta-training and meta-testing tasks on base and novel fault categories of the involved system(s), as described below:
@@
In the meta-testing phase of each experiment, 50 distinct meta-testing tasks are randomly created by selecting 5 distinct fault categories out of 10 novel fault categories of the involved system. The mathematical permutations for selecting 5 out of 10 categories yields a total of 252 possible permutations, calculated as $C(10,5)$. We focus on 50 distinct meta-testing tasks that represent 252 possible permutations to evaluate our framework following the prior studies [ye2021train, finn2017model]. This approach allows for a focused comprehensive evaluation, ensuring a broad coverage of various fault categories while keeping the data size within practical limits for detailed examination. Each meta-testing task is unique and is either a 5-way 5-shot or 10-shot abnormal trace classification task.
@@ IV-A4 Baselines
As the MultiHAttenAE alternative, OnlySpan follows the related work [nedelkoski2019anomaly2] to consider each trace as a sequence of spans and construct trace representations using span attributes. As alternatives to the multi-head attention fusion (used in our MultiHAttenAE), Linear-AE and GLU-AE use the linear projection and gate mechanism, respectively, to fuse spans and logs for generating trace representations. These approaches provide simpler alternatives to multi-head attention mechanisms. Meta-learner alternatives include the basic linear model, sequence models RNN and LSTM, and CNN. MAML alternatives include metric-based meta-learning algorithms (ProtoNet [snell2017prototypical], MatchingNet [vinyals2016matching]) and traditional classification models (NearNeighbor and Decision tree). These alternatives are simpler than MAML. Note that, MAML alternative baselines are only evaluated on E1 and E3 since the fundamental algorithms of these do not have cross-system adaptability.
@@ IV-B Evaluation results
Model E1. TrainTicket to TrainTicket E3. OnlineBoutique to TrainTicket
5-shot 10-shot 5-shot 10-shot
Our MultiHAttenAE+TE-MAML 92.91$\pm$2.10 (74.67-100.0) 93.26$\pm$1.40 (76.00-100.0) 86.35$\pm$2.00 (70.67-100.0) 92.19$\pm$1.99 (74.67-100.0)
@@
Linear-AE+TE-MAML 89.15$\pm$2.29 (73.33-100.0) 90.59$\pm$2.43 (70.67-100.0) 83.09$\pm$2.55 (62.67-97.33) 90.61$\pm$2.01 (72.00-100.0)
GLU-AE +TE-MAML 92.21$\pm$1.73 (77.33-100.0) 93.07$\pm$1.64 (77.33-100.0) 85.07$\pm$2.38 (66.67-100.0) 94.40$\pm$2.19 (72.00-100.0)
@@
MultiHAttenAE+Linear-MAML 45.84$\pm$2.21 (25.33-61.33) 45.36$\pm$2.16 (30.67-60.00) 43.81$\pm$1.99 (28.00-58.67) 43.87$\pm$1.93 (29.33-60.00)
MultiHAttenAE+RNN-MAML 49.65$\pm$2.09 (37.33-64.00) 42.88$\pm$1.93 (24.00-58.67) 48.45$\pm$1.75 (38.67-65.33) 47.07$\pm$1.88 (34.67-58.67)
MultiHAttenAE+LSTM-MAML 41.39$\pm$2.20 (21.33-56.00) 42.67$\pm$1.91 (29.33-56.00) 40.32$\pm$1.68 (22.67-52.00) 42.29$\pm$1.79 (25.33-56.00)
MultiHAttenAE+CNN-MAML 57.06$\pm$2.85 (41.33-81.00) 69.20$\pm$2.19 (56.00-88.00) 49.04$\pm$1.80 (38.67-64.00) 69.47$\pm$2.65 (48.00-89.33)
@@
MultiHAttenAE+TE-MatchingNet 76.56$\pm$2.80 (49.33-93.33) 76.05$\pm$2.37 (50.67-94.67)
MultiHAttenAE+ProtoNet 57.25$\pm$0.03 (40.00-74.67) 59.68$\pm$0.03 (44.00-76.00)
MultiHAttenAE+NearNeighbor 88.19$\pm$0.02 (74.67-98.67) 92.56$\pm$0.02 (78.00-100.0)
MultiHAttenAE+DecisonTree 66.80$\pm$0.03 (46.67-88.00) 77.09$\pm$0.03 (54.67-96.00)
@@
Model E2. OnlineBoutique to OnlineBoutique E4. TrainTicket to OnlineBoutique
5-shot 10-shot 5-shot 10-shot
Our MultiHAttenAE+TE-MAML 82.50$\pm$2.35 (65.33-98.67) 85.20$\pm$2.33 (66.67-98.67) 82.37$\pm$2.07 (64.00-97.33) 84.77$\pm$2.28 (68.00-98.67)
@@
Linear-AE+TE-MAML 76.15$\pm$2.59 (60.00-95.00) 78.21$\pm$2.50 (64.00-96.00) 75.81$\pm$2.45 (52.00-89.33) 74.32$\pm$2.44 (53.33-88.00)
GLU-AE +TE-MAML 80.61$\pm$2.96 (58.67-98.67) 77.49$\pm$2.67 (48.00-94.67) 74.96$\pm$2.76 (54.67-94.67) 77.57$\pm$2.7 (56.00-94.67)
@@
MultiHAttenAE+Linear-MAML 42.59$\pm$3.63 (20.00-77.33) 40.75$\pm$3.53 (21.33-68.00) 47.01$\pm$3.59 (25.33-74.67) 44.35$\pm$4.10 (20.00-89.33)
MultiHAttenAE+RNN-MAML 72.59$\pm$2.50 (54.67-94.67) 64.75$\pm$2.53 (46.67-80.00) 72.58$\pm$2.58 (54.70-94.70) 71.01$\pm$2.80 (56.00-89.33)
MultiHAttenAE+LSTM-MAML 54.80$\pm$2.11 (40.00-70.67) 55.97$\pm$2.25 (41.33-69.33) 56.19$\pm$1.92 (38.67-72.00) 59.71$\pm$2.05 (42.67-77.33)
MultiHAttenAE+CNN-MAML 80.10$\pm$2.16 (60.00-94.67) 83.07$\pm$3.29 (68.00-97.33) 79.01$\pm$2.63 (56.00-97.33) 84.08$\pm$2.76 (65.33-100.0)
@@
MultiHAttenAE+TE-MatchingNet 76.29$\pm$3.00 (54.67-96.00) 73.11$\pm$2.94 (50.67-94.67)
MultiHAttenAE+ProtoNet 74.51$\pm$0.03 (53.33-92.00) 76.59$\pm$0.04 (58.66-94.67)
MultiHAttenAE+NearNeighbor 80.96$\pm$0.03 (64.00-98.67) 84.75$\pm$0.03 (62.80-98.67)
MultiHAttenAE+DecisonTree 66.99$\pm$0.02 (54.67-80.00) 73.79$\pm$0.03 (58.67-82.67)
@@
5-shot 10-shot 5-shot 10-shot
Our framework 0.0460 0.0680 0.0651 0.0953
GLU-AE +TE-MAML 0.0640 0.0942 0.0651 0.0954
Linear-AE +TE-MAML 0.0748 0.0946 0.0640 0.0944
MultiHAttenAE+NearNeighbor 0.1860 0.1890
@@
5-shot 10-shot 5-shot 10-shot
Our framework 0.0848 0.0977 0.0875 0.0974
GLU-AE +TE-MAML 0.0801 0.0978 0.0731 0.0969
CNN-MAML 0.0470 0.0472 0.0417 0.0490
MultiHAttenAE+NearNeighbor 0.5140 0.5307
@@
GLU-AE+TE-MAML 1.8395 2.1541 4.3121 5.1921
Linear-AE+TE-MAML 1.8286 2.1486 4.1223 5.2066
OnlySpan+TE-MAML 1.1671 1.1818 1.9065 2.1822
@@ V Threats to validity
The main internal threats may lay in the correctness of baselines’ development. Given that there is only one related work [nedelkoski2019anomaly2], we designed most of the baselines ourselves based on our understanding via an ablation study. To mitigate this threat, we carefully reviewed the core principles of the baseline methodologies, conducted extensive testing, and consulted domain experts for their insights.
@@
The external threats may arise from the limitation of our fault dataset sourced from open datasets DeepTraLog and Nezha. First, we intended to include trace-related metrics in constructing the latent trace representations for MSS. Including trace-related metrics could potentially help in distinguishing resource-related anomalies within a trace. This may address the issue referred to RQ1 results, where our framework and baselines are less effective in differentiating abnormal traces caused by CPU contention and network delay in OnlineBoutique, as well as those caused by network delay and service latency in Trainticket. Nezha is a dataset that contains sufficient modality data and enables the extraction of spans, logs, and metrics within a trace to construct its representation for the observed MSS. However, we did not find other similar MSS open datasets that meet such needs. Second, we aimed to expand the evaluation of our framework by including additional datasets that contain abnormal traces from a broader range of fault categories, as well as datasets from various MSS. Despite our efforts, we were unable to find such datasets. Thus, our study used abnormal traces from 30 and 32 fault categories in Trainticket and OnlineBoutique, respectively. This limitation led us to adopt a 5-way setup in order to ensure an adequate distribution of meta-training and meta-testing tasks in each experiment. However, prior studies have demonstrated that MAML is highly effective for complex text classification tasks (e.g., 20-way and 50-way classifications) across a variety of contexts [jung2022few, jamal2019task, yu2020transmatch]. The above two external threats pose the possibility that our application of our framework for classifying abnormal traces for MSS might not fully utilize its generalization capabilities. Addressing these threats may improve our framework, thereby enhancing the precision and robustness of trace-level RCA in MSS. To address these threats, we are in the process of deploying MSS and generating new open datasets.
VI Related work
{This part is still under writing}
Research interest in fault case classification and prediction for MSS and similar distributed systems has significantly grown. Existing machine learning methods can be divided into two groups: label classification-based and large language model-based (LLM-based).
(1) Label classification-based: These methods involve training models to classify or predict faults based on predefined labels. Many MSS-related studies use binary fault classification models to predict system behaviors using two labels - normal (non-faulty) and abnormal (faulty)- with data traces, logs, and metrics. Besides, two studies [zhou2019latent, nedelkoski2019anomaly2] built multi-label classification models to predict trace-based fault categories. The details are described below.
The study [zhou2019latent] trained a multi-label classification model using an extensive collection of trace logs (i.e., system logs) on two benchmark MSS, Sock Shop and Trainticket. The model’s performance was evaluated on Trainticket across 10 fault categories, achieving a recall 0.647-0.983 (0.821 on average) and a precision 0.586-0.984 (0.891 on average). The prediction accuracy on some faults like configuration-related faults is lower than others.
The study [nedelkoski2019anomaly2] uses a convolutional neural network (CNN) model to classify trace-related anomalies (based on their patterns in response time series) into 15 pre-defined categories, e.g., incremental, mean shift, gradual increase, cylinder. It trained and tested the model using a large dataset from a real-world production cloud. In the testing phase, the model accurately classifies the tested anomaly patterns, and achieved 80% and 48% accuracy when Gaussian noise was added with a = [0.05 and 0.1] respectively.
(3) LLM-based: These studies [gupta2023learning, chen2023empowering, zhang2024automated] leverage the power of LLMs to classify fault categories and analyze root cause with few labeled data. The details are described below.
The study [gupta2023learning] built a Transformer-based LLM for AIOps to classify log lines or messages into seven fault categories: Memory, Network, Authentication, I/O, Device, Application, and Other. It involved 7 experts in labeling logs, and fine-tuned the model using 10, 20, and 30 shots for each fault category, which come from a total pool of 868 labeled data across fault categories. In testing with the remaining labeled data, the 30-shot setting was the most effective, the model accurately predicted Memory and I/O related faults, while achieving 69.54-76.24% accuracy for the other four fault categories.
The study [chen2023empowering] proposed a LLM-based system for analyzing root causes of cloud incidents, using traces, logs, and metrics. The system is designed to assign incoming incidents to appropriate handlers and to categorize them into 10 categories: authorization, service running, configuration, I/O, invalid value, network, attack, code bug, memory, and service invocation. This study collected a dataset of 653 incidents from Microsoft’s Transport, each incident is labeled by experts with the specific fault category. It uses 75% of the dataset to train the system. The system achieved a micro F1-score of 0.766 and a macro F1-score of 0.533 in predicting the fault category for cloud incidents. This indicates the system is generally reliable at predicting individual instances across all fault categories, but the lower macro F1-score suggests that the system may struggle with some fault categories.
Research interest in fault case classification and prediction for MSS and comparable cloud-based infrastructures has significantly grown. However, the adoption of machine learning techniques in this domain was hindered by a lack of sufficiently labeled observational data, including traces, logs, and metrics, as well as by the heterogeneity of these data types. Both factors are essential for the effective deployment of machine learning algorithms

### v3

@@ III-C2 Meta-training
The standard MAML outer loop update involves a computationally intensive process of computing a gradient through a gradient, which requests an additional backward pass through the function $f$ to compute Hessian-vector products [18]. We use a first-order approximation [18] to simplify the update by such second-order derivatives.
@@ IV-A1 Dataset
We use two open-source datasets to establish our own dataset: DeepTraLog [32] and Nezha [33]. (1) DeepTraLog was collected from TrainTicket, a medium-scale, widely-used benchmark MSS for train ticket booking, comprising 45 services. TrainTicket contains various real-industrial faults in different branches. DeepTraLog was created by Zhang et al. [9], who executed such fault branches of TrainTicket and generated abnormal traces of 14 fault categories. These fault categories pertain to related to asynchronous interaction, multi-instance, configuration, and monolithic dimensions. (2) Nezha was created by Yu et al. [34]. It includes data from two MSS: TrainTicket, and a small-scale e-commerce benchmark MSS OnlineBoutique composed of 12 services. This dataset was constructed by injecting faults into the pods of TrainTicket and OnlineBoutique and collocating abnormal traces. Within each system, faults injected into each pod belong to four categories: CPU-intention, CPU-consumption, service exception, and message return. Each fault case within a pod here is considered as a unique fault category.
@@ IV-A2 Training and evaluation
E1 and E3 are within system experiments while E2 and E4 are cross-system experiments. Considering prior MAML studies [35, 18] and our fault dataset, we investigate 5-way 5-shot and 10-shot setups across experiments E1-E4. In these experiments, each meta-training and meta-testing task is a unique abnormal trace classification task, where each task involves classifying abnormal traces from a distinct set of 5-way fault categories. In each experiment, we train and evaluate our TE-MAML using the randomly established meta-training and meta-testing tasks on base and novel fault categories of the involved system(s), as described below:
@@
In the meta-testing phase of each experiment, 50 distinct meta-testing tasks are randomly created by selecting 5 distinct fault categories out of 10 novel fault categories of the involved system. The mathematical permutations for selecting 5 out of 10 categories yields a total of 252 possible permutations, calculated as $C(10,5)$. We focus on 50 distinct meta-testing tasks that represent 252 possible permutations to evaluate our framework following the prior studies [35, 18]. This approach allows for a focused comprehensive evaluation, ensuring a broad coverage of various fault categories while keeping the data size within practical limits for detailed examination. Each meta-testing task is unique and is either a 5-way 5-shot or 10-shot abnormal trace classification task.
@@ IV-A4 Baselines
As the MultiHAttenAE alternative, OnlySpan follows the related work [23] to consider each trace as a sequence of spans and construct trace representations using span attributes. As alternatives to the multi-head attention fusion (used in our MultiHAttenAE), Linear-AE and GLU-AE use the linear projection and gate mechanism, respectively, to fuse spans and logs for generating trace representations. These approaches provide simpler alternatives to multi-head attention mechanisms. Meta-learner alternatives include the basic linear model, sequence models RNN and LSTM, and CNN. MAML alternatives include metric-based meta-learning algorithms (ProtoNet [36], MatchingNet [37]) and traditional classification models (NearNeighbor and Decision tree). These alternatives are simpler than MAML. Note that, MAML alternative baselines are only evaluated on E1 and E3 since the fundamental algorithms of these do not have cross-system adaptability.
@@ IV-B Evaluation results
Model
E1. TrainTicket to TrainTicket E3. OnlineBoutique to TrainTicket
5-shot
10-shot
5-shot
10-shot
Our MultiHAttenAE+TE-MAML
92.91$\pm$2.10 (74.67-100.0)
93.26$\pm$1.40 (76.00-100.0)
86.35$\pm$2.00 (70.67-100.0)
92.19$\pm$1.99 (74.67-100.0)
@@
Linear-AE+TE-MAML
89.15$\pm$2.29 (73.33-100.0)
90.59$\pm$2.43 (70.67-100.0)
83.09$\pm$2.55 (62.67-97.33)
90.61$\pm$2.01 (72.00-100.0)
GLU-AE +TE-MAML
92.21$\pm$1.73 (77.33-100.0)
93.07$\pm$1.64 (77.33-100.0)
85.07$\pm$2.38 (66.67-100.0)
94.40$\pm$2.19 (72.00-100.0)
@@
MultiHAttenAE+Linear-MAML
45.84$\pm$2.21 (25.33-61.33)
45.36$\pm$2.16 (30.67-60.00)
43.81$\pm$1.99 (28.00-58.67)
43.87$\pm$1.93 (29.33-60.00)
MultiHAttenAE+RNN-MAML
49.65$\pm$2.09 (37.33-64.00)
42.88$\pm$1.93 (24.00-58.67)
48.45$\pm$1.75 (38.67-65.33)
47.07$\pm$1.88 (34.67-58.67)
MultiHAttenAE+LSTM-MAML
41.39$\pm$2.20 (21.33-56.00)
42.67$\pm$1.91 (29.33-56.00)
40.32$\pm$1.68 (22.67-52.00)
42.29$\pm$1.79 (25.33-56.00)
MultiHAttenAE+CNN-MAML
57.06$\pm$2.85 (41.33-81.00)
69.20$\pm$2.19 (56.00-88.00)
49.04$\pm$1.80 (38.67-64.00)
69.47$\pm$2.65 (48.00-89.33)
@@
MultiHAttenAE+TE-MatchingNet
76.56$\pm$2.80 (49.33-93.33)
76.05$\pm$2.37 (50.67-94.67)
MultiHAttenAE+ProtoNet
57.25$\pm$0.03 (40.00-74.67)
59.68$\pm$0.03 (44.00-76.00)
MultiHAttenAE+NearNeighbor
88.19$\pm$0.02 (74.67-98.67)
92.56$\pm$0.02 (78.00-100.0)
MultiHAttenAE+DecisonTree
66.80$\pm$0.03 (46.67-88.00)
77.09$\pm$0.03 (54.67-96.00)
@@
Model
E2. OnlineBoutique to OnlineBoutique E4. TrainTicket to OnlineBoutique
5-shot
10-shot
5-shot
10-shot
Our MultiHAttenAE+TE-MAML
82.50$\pm$2.35 (65.33-98.67)
85.20$\pm$2.33 (66.67-98.67)
82.37$\pm$2.07 (64.00-97.33)
84.77$\pm$2.28 (68.00-98.67)
@@
Linear-AE+TE-MAML
76.15$\pm$2.59 (60.00-95.00)
78.21$\pm$2.50 (64.00-96.00)
75.81$\pm$2.45 (52.00-89.33)
74.32$\pm$2.44 (53.33-88.00)
GLU-AE +TE-MAML
80.61$\pm$2.96 (58.67-98.67)
77.49$\pm$2.67 (48.00-94.67)
74.96$\pm$2.76 (54.67-94.67)
77.57$\pm$2.7 (56.00-94.67)
@@
MultiHAttenAE+Linear-MAML
42.59$\pm$3.63 (20.00-77.33)
40.75$\pm$3.53 (21.33-68.00)
47.01$\pm$3.59 (25.33-74.67)
44.35$\pm$4.10 (20.00-89.33)
MultiHAttenAE+RNN-MAML
72.59$\pm$2.50 (54.67-94.67)
64.75$\pm$2.53 (46.67-80.00)
72.58$\pm$2.58 (54.70-94.70)
71.01$\pm$2.80 (56.00-89.33)
MultiHAttenAE+LSTM-MAML
54.80$\pm$2.11 (40.00-70.67)
55.97$\pm$2.25 (41.33-69.33)
56.19$\pm$1.92 (38.67-72.00)
59.71$\pm$2.05 (42.67-77.33)
MultiHAttenAE+CNN-MAML
80.10$\pm$2.16 (60.00-94.67)
83.07$\pm$3.29 (68.00-97.33)
79.01$\pm$2.63 (56.00-97.33)
84.08$\pm$2.76 (65.33-100.0)
@@
MultiHAttenAE+TE-MatchingNet
76.29$\pm$3.00 (54.67-96.00)
73.11$\pm$2.94 (50.67-94.67)
MultiHAttenAE+ProtoNet
74.51$\pm$0.03 (53.33-92.00)
76.59$\pm$0.04 (58.66-94.67)
MultiHAttenAE+NearNeighbor
80.96$\pm$0.03 (64.00-98.67)
84.75$\pm$0.03 (62.80-98.67)
MultiHAttenAE+DecisonTree
66.99$\pm$0.02 (54.67-80.00)
73.79$\pm$0.03 (58.67-82.67)
@@
5-shot
10-shot
5-shot
10-shot
Our framework
0.0460
0.0680
0.0651
0.0953
GLU-AE +TE-MAML
0.0640
0.0942
0.0651
0.0954
Linear-AE +TE-MAML
0.0748
0.0946
0.0640
0.0944
MultiHAttenAE+NearNeighbor
0.1860
0.1890
@@
5-shot
10-shot
5-shot
10-shot
Our framework
0.0848
0.0977
0.0875
0.0974
GLU-AE +TE-MAML
0.0801
0.0978
0.0731
0.0969
CNN-MAML
0.0470
0.0472
0.0417
0.0490
MultiHAttenAE+NearNeighbor
0.5140
0.5307
@@
GLU-AE+TE-MAML
1.8395 2.1541 4.3121 5.1921
Linear-AE+TE-MAML
1.8286 2.1486 4.1223 5.2066
OnlySpan+TE-MAML
1.1671 1.1818 1.9065 2.1822
@@ V Threats to validity
The main internal threats may lay in the correctness of baselines’ development. Given that there is only one related work [23], we designed most of the baselines ourselves based on our understanding via an ablation study. To mitigate this threat, we carefully reviewed the core principles of the baseline methodologies, conducted extensive testing, and consulted domain experts for their insights.
@@
The external threats may arise from the limitation of our fault dataset sourced from open datasets DeepTraLog and Nezha. First, we intended to include trace-related metrics in constructing the latent trace representations for MSS. Including trace-related metrics could potentially help in distinguishing resource-related anomalies within a trace. This may address the issue referred to RQ1 results, where our framework and baselines are less effective in differentiating abnormal traces caused by CPU contention and network delay in OnlineBoutique, as well as those caused by network delay and service latency in Trainticket. Nezha is a dataset that contains sufficient modality data and enables the extraction of spans, logs, and metrics within a trace to construct its representation for the observed MSS. However, we did not find other similar MSS open datasets that meet such needs. Second, we aimed to expand the evaluation of our framework by including additional datasets that contain abnormal traces from a broader range of fault categories, as well as datasets from various MSS. Despite our efforts, we were unable to find such datasets. Thus, our study used abnormal traces from 30 and 32 fault categories in Trainticket and OnlineBoutique, respectively. This limitation led us to adopt a 5-way setup in order to ensure an adequate distribution of meta-training and meta-testing tasks in each experiment. However, prior studies have demonstrated that MAML is highly effective for complex text classification tasks (e.g., 20-way and 50-way classifications) across a variety of contexts [38, 39, 40]. The above two external threats pose the possibility that our application of our framework for classifying abnormal traces for MSS might not fully utilize its generalization capabilities. Addressing these threats may improve our framework, thereby enhancing the precision and robustness of trace-level RCA in MSS. To address these threats, we are in the process of deploying MSS and generating new open datasets.
VI Conclusion
This paper proposes a framework for few-shot abnormal trace classification for MSS with two main components: (1) MultiHAttenAE for constructing system-specific trace representations, which enables (2) TE-MAML to perform effective and efficient few-shot learning for abnormal trace classification. The proposed framework is evaluated on representative benchmark MSS with open datasets. The evaluation results shows our framework’s effective and efficient within-system and cross-system adaptability. Overall, our framework provides a solution for a few-shot abnormal trace classification for MSS. Future work will focus on further improving the generalizability, scalability, and interpretability of the proposed framework. Also, we plan to evaluate its performance on real industrial MSS.
VII Replication package
To ensure the transparency and reproducibility of our research, we will publish a replication package upon acceptance of the paper. The replication package will include all the necessary artifacts to replicate our experiments, such as the source code, data, and implementation details. We will provide detailed instructions on how to use the replication package and will make it available on a GitHub public repository.
VIII Acknowledgment
The authors would like to thank Nauman Ali for reading the paper and providing valuable feedback.